If x and y and are positive numbers connected by the relation y = ax²/(x-a) (x-b) (x-c) + bx/(x-b) (x-c) + c / (x-c) + 1 then prove that log y = 3log x-log | x-a|- log | x-b|- log | x-c|
for any valid base of logarithms.

Grade:

1 Answers

Prudhvi teja
13 years ago

Daer ajit

simplify the equation

y = ax2+ bx(x-a)+c(x-a)(x-c)/(x-a)(x-b)x-c) +1

ax2+ bx(x-a)+c(x-a)(x-c =x3- (x-a)(x-b)(x-c)+1

y=x3/(x-a)(x-b)(x-c) -1+1

so

log(y) = 3log x-log | x-a|- log | x-b|- log | x-c|

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly.

All the best.

Regards,
Askiitians Experts
Prudhvi Teja

 

 


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian